HUBBLE SPACE TELESCOPE DAILY REPORT #826 PERIOD COVERED: 0800 EST Daily Status Report as of 049/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 3483 (Lyman Alpha Emission in Galaxies) The Faint Object Spectrograph was used on both the red and blue sides to make six observations looking for Lyman Alpha emissions from a near by galaxy which lies near the line of sight of quasi-stellar object Q1209+968. This will help confirm that star formation is taking place. The first five observations were received and displayed. They were reported to look noisy, but may look better after being co-added. The last observations was delayed due to the data line between GSFC and STScI being down for 3.5 hours. After the line was fixed, the image was reported to look okay. 1.2 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make observations of the protoplanetary system PK349+01D1 to study its current morphology. The observations were executed as scheduled and no problems were reported. 1.3 Completed PC 3799 (PC Imaging of Gravitation Lenses, WFPC GTO Augmentation) The Planetary camera was used to take three observations of the gravitational lens system PG1115+080 to obtain data to aid in the characterization of the lensing system. The three observations were executed on Single Star Course Track using FGS #1. The first two observations were received and displayed. They were reported to look good, and there was no measurable motion in the target. A second gyro bias update was uplinked before the final observation. No problems were reported. 1.4 Completed FGS 1570 (Extragalactic Astrometry and Astrophysics F10-11, PT 10F2) The FGSs were used to attempt astrometric observations of extra galactic source 0818-128IN. The astrometric observations were supposed to be on FGS #3, using FGS #1 for evidence. Instead, the high voltage was on for FGS #2, and off for FGS #3. FGS #2 servos were not spinning. HSTAR 3888 was written. (See Note 4.2) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: NOTE: Only FGS #1 and #3 are being used. Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 7 Successful : 7 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Significant HSTARs: None 2.4 Operations Note: The GSFC Simulation #4 (HST Deploy) was completed yesterday (048/1300Z - 049/0100Z). All test objectives were met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of PC 3603 (Snapshots of Protoplanetary Nebulae) 3.2 Complete GHRS/2 3862 (Excitation of Atmospheres of Planetary Satellites) 3.3 Complete PC 4140 (WF/PC Cal: Rapid Internal Monitor- Part 2) 3.4 Complete GHRS/2 4065 (Dark Noise Monitoring HRS) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.6 Complete GHRS/2 4792 (Hydrogen Lyman Alpha Emission Line Profiles) 3.7 Initiate GHRS/2 3372 (GHRS Point Spread Function Determination) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: The FGS #2 anomalous behavior is currently under study. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END- /DMH